OpenStudy (anonymous):

The sum of three consecutive integers is -78. Determine the smallest integer.

mathslover (mathslover):

let the integers be x-1 , x , x+1 sum = \[x-1+x+x+1 = -78\] \[3x=-78\] \[x=-78/3\] \[x=-26\] smallest ineger : x-1 = -26-1 = -27

OpenStudy (anonymous):

3x+3=-78,x+1=-26,x=-27

OpenStudy (anonymous):

let the smallest integer be 'x'. Then three integers are x, x+1, and x+2 now their sum = x + (x+1) + (x+2) = -78 Or, 3x + 3 = -78 or, x = -27 = smallest integer.

mathslover (mathslover):

Remember in these cases : u must take the consecutive integers as; x-1 , x , x + 1 as the calculation becomes easy ( -1 and1 gets cut down )